what is the slope of the line 5x+2y=-2
Found 2 solutions by rfer, lenny460:
Answer by rfer(16322)

2y=-5x-2
y=-2.5x-1
slope m=-2.5

Answer by lenny460(1073)

Plot the straight line graph:
5x + 2y = -2
Subtract 5x from each side of the equation.
5x + 2y - 5x = -2 - 5x
Combine like terms:
5x - 5x = 0
We now have:
2y = -2 - 5x or
2y = -5x - 2
Divide each side by 2
2y/2 = -5x/2 - 2/2
y = -5x/2 - 1

From the above we get the following x and y coordinates:
(-4,9)
(-2,4)
(0,-1)
(2,-6)
(4,-11)

The Slope Formula:
(y2 - y1)/(x2 - x1)
(-6 - (-1)/(2 - 0)
-6 - (-1) = -6 + 1 = -5
2 - 0 = 2


The Slope (m) = -5/2 (or -2.5)



obuong3@aol.com